# TU Munich, Germany

### v1

@@ 1 Introduction
Medical ultrasound (US) is widely used in modern clinical practice for examining internal organs such as the carotid, thyroid, and liver. With its accessibility and portability, US imaging has the potential for widespread deployment, making it more universally available. However, interpreting US images requires substantial experience due to significant visual variability stemming from differences in imaging and acquisition parameters. Unlike CT and MRI, US interpretation is less intuitive due to its limited field of view and lack of 3D structural information. Therefore, effective US image explanation and scanning guidance are essential for broader adoption, particularly in point-of-care ultrasound (POCUS) settings [22, 2] and for self-learning of anatomy and physiology among non-experts.
@@
To demonstrate the impact of SG representation in US imaging, this study introduces a novel method that leverages highly conceptualized SGs for realizing two critical tasks essential to advancing the democratization of portable and accessible US imaging: (1) generating graspable US explanation for ordinaries to self-learn anatomical and physiological knowledge, which is also particularly useful for POCUS scenarios, and (2) providing scanning guidance to reveal missing anatomies in the current imaging view, ensuring the displayed content aligns with user preferences. Both US summary [9, 14] and probe guidance [30, 19, 28] tasks are important and have been investigated. However, to the best of our knowledge, this is the first work introducing SG and LLMs to boost intuitive US explanation and scanning guidance.
@@ 3.1 Implementation Details
Model Selection We use hyperparameters similar to those in RelTR [6] for our experiments. The SG prediction network is trained for 800 epochs with a batch size of 16, on a workstation equipped with an RTX 4080 Super GPU. The initial learning rate for the transformer is $10^{-4}$ and is reduced by a factor of 0.1 after 200 epochs to ensure stable convergence. Since this work is aimed at potential applications in portable US devices, we prioritize “lightweight" LLMs like LLaMa [27], Qwen [3], Gemma [26], Mathstral [21] and the distilled DeepSeek R1 using the Qwen model (DS-R1-Qwen) [8]. These lightweight models are quantized versions optimized for efficiency, reducing computational demands while maintaining performance, making them more suitable for real-time processing in resource-constrained environments. Additionally, we also employ high-capacity LLMs, such as Gemini 2.0 Flash [25], and Grok 3 as reference models.
Data Acquisition The carotid artery US images were acquired using Siemens Juniper US Machine (ACUSON Juniper, SIEMENS AG, Germany) equipped with a 12L3 linear probe. The imaging and focus depths were set to 45 mm and 20 mm, respectively. A total of 289 US images were collected and annotated, with the training set comprising 262 images (resolution: 829$\times$770 pixels) from five volunteers. An additional 27 images collected from different volunteers were used for testing. Due to the lack of mature tools for scene graph annotation, we developed a lightweight annotation tool specifically for our 2D US dataset [^1] The annotation tool and dataset will be released upon acceptance.. It worth noting that, in addition to object detection annotations, scene graph annotation requires labeling triplets (<$subject-predicate-object$>), which makes the annotation process significantly more labor-intensive. Moreover, in SG prediction tasks, the predicate in each triplet is closely tied to the spatial relationships between entities in the image. As a result, standard data augmentation techniques used in traditional computer vision tasks are mostly inapplicable, with horizontal flipping being one of the few exceptions. These challenges have led to a relatively small dataset size.
@@ 4 Discussion and Conclusion
4.0.1 Acknowledgements
This work was supported in part by the Multi-Scale Medical Robotics Center, AIR@InnoHK, Hong Kong; and in part by the SINO-German Mobility Project under Grant M0221.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ 1 Introduction
Medical ultrasound (US) is widely used in modern clinical practice for examining internal organs such as the carotid, thyroid, and liver. With its accessibility and portability, US imaging has the potential for widespread deployment, making it more universally available. However, interpreting US images requires substantial experience due to significant visual variability stemming from differences in imaging and acquisition parameters. Unlike CT and MRI, US interpretation is less intuitive due to its limited field of view and lack of 3D structural information. Therefore, effective US image explanation and scanning guidance are essential for broader adoption, particularly in point-of-care ultrasound (POCUS) settings [24, 2, 16, 5] and for self-learning of anatomy and physiology among non-experts.
@@
To demonstrate the impact of SG representation in US imaging, this study introduces a novel method that leverages highly conceptualized SGs for realizing two critical tasks essential to advancing the democratization of portable and accessible US imaging: (1) generating graspable US explanation for ordinaries to self-learn anatomical and physiological knowledge, which is also particularly useful for POCUS scenarios, and (2) providing scanning guidance to reveal missing anatomies in the current imaging view, ensuring the displayed content aligns with user preferences. Both US summary [10, 17, 33] and probe guidance [32, 22, 30, 15] tasks are important and have been investigated. However, to the best of our knowledge, this is the first work introducing SG and LLMs to boost intuitive US explanation and scanning guidance.
@@ 3.1 Implementation Details
Model Selection We use hyperparameters similar to those in RelTR [7] for our experiments. The SG prediction network is trained for 800 epochs with a batch size of 16, on a workstation equipped with an RTX 4080 Super GPU. The initial learning rate for the transformer is $10^{-4}$ and is reduced by a factor of 0.1 after 200 epochs to ensure stable convergence. Since this work is aimed at potential applications in portable US devices, we prioritize “lightweight" LLMs like LLaMa [29], Qwen [3], Gemma [28], Mathstral [^1] https://mistral.ai/ and the distilled DeepSeek R1 using the Qwen model (DS-R1-Qwen) [9]. These lightweight models are quantized versions optimized for efficiency, reducing computational demands while maintaining performance, making them more suitable for real-time processing in resource-constrained environments. Additionally, we also employ high-capacity LLMs, such as Gemini 2.0 Flash [27], and Grok 3 as reference models.
Data Acquisition The carotid artery US images were acquired using Siemens Juniper US Machine (ACUSON Juniper, SIEMENS AG, Germany) equipped with a 12L3 linear probe. The imaging and focus depths were set to 45 mm and 20 mm, respectively. A total of 289 US images were collected and annotated, with the training set comprising 262 images (resolution: 829$\times$770 pixels) from five volunteers. An additional 27 images collected from different volunteers were used for testing. Due to the lack of mature tools for scene graph annotation, we developed a lightweight annotation tool specifically for our 2D US dataset. It worth noting that, in addition to object detection annotations, scene graph annotation requires labeling triplets (<$subject-predicate-object$>), which makes the annotation process significantly more labor-intensive. Moreover, in SG prediction tasks, the predicate in each triplet is closely tied to the spatial relationships between entities in the image. As a result, standard data augmentation techniques used in traditional computer vision tasks are mostly inapplicable, with horizontal flipping being one of the few exceptions. These challenges have led to a relatively small dataset size.